### v2

@@ III-B Target Sensing Only
CRB is the lower bound of the unbiased estimation variance, and therefore minimizing CRB is a strategy for improving the accuracy of the estimation555 There are three different types of scalar CRB metrics based on the trace, maximum eigenvalue, and determinant of CRB matrix [12, 38]. In this work, the determinant of the CRB matrix is chosen as the performance metric of TS. The proposed schemes in this work also fit those scenarios wherein the trace and the maximum eigenvalue of the CRB are utilized as the matrices. . In this subsection, the transmitting beam is designed only to minimize the determinant of CRB.
@@ III-B2 Imperfect Target Direction
In this subsection, more practical scenario is considered wherein the estimated target angle ${\theta_{T_{k}}}$ is imperfect and there is an uncertain angle $\Delta\theta$. Specifically, the angle of the $k$-th target varies in ${\Omega_{1}}=\left[{{\theta_{T_{k}}}-\Delta\theta,{\theta_{T_{k}}}+\Delta\theta}\right]$. Then, the following optimization problem is formulated
@@ IV Jointly Optimization of Proactive Eavesdropping and TS
where $\alpha_{\Phi}$ denotes the target threshold for determinant of the CRB.
@@ IV-A Proactive Eavesdropping-centric Design
where (20e) is the TS constraint and $\alpha_{\Phi}$ denotes the target threshold for determinant of the CRB. $\mathcal{P}_{3}$ is a convex problem after dropping constraint (20c), which can be solved by CVX toolbox.
@@ IV-B Target Sensing-centric Design
where $\alpha_{I}$ denotes the interference threshold at $D$. It must be noted that (21f) and (22f) denote the requirement for the eavesdropping rate. Obviously, $\alpha_{I}$ has an essential effect on the optimized results. $\mathcal{P}_{4.1}$ and $\mathcal{P}_{4.2}$ can be transformed into a convex problem by means of SDR and are solved directly using the CVX toolbox when (21c) and (22c) are omitted.
@@
In this subsection, a normalized weighted optimization problem that reveals the performance trade-off between PE and TS is proposed. Design of the weighted optimization between the CRB and eavesdropping rate presents the challenge that both two performance metrics have different units and potentially magnitudes. Inspired by [24], both $\mathcal{I}\left({{\theta_{D}}}\right)$ and ${\left|{{\bm{\Phi}}}\left({{\bm{\zeta}}}\right)\right|}$ are normalized with their respective bound obtained in $\mathcal{P}_{1.1}$ and $\mathcal{P}_{2.1}$, respectively.
@@ IV-C1 Joint Optimization Proactive Eavesdropping and Target Sensing
To achieve a trade-off between PE and TS, while taking the uncertain angle of $T_{k}$ into account, the weighted optimization problem is formulated as
@@
[…]
@@ IV-C2 The interference power is zero
It is unnecessary for $E$ to send AN to $D$ when the channel quality of $S-E$ is better than that of $S-D$, i.e. $\frac{{h_{SD}^{2}}}{{\sigma_{D}^{2}}}\leq\frac{{h_{SE}^{2}}}{{\sigma_{E}^{2}}}$ in (23d) [41]. In these scenarios, the eavesdropping rate is equal to the receiving rate of $D$ without interference and $\mathcal{P}_{3}$ is transformed into a TS sub-problem, which is expressed as
@@
       Solve $\mathcal{P}_{1.1}$ for given parameters and obtain the solution ~${{{\mathrm{\tilde{\mathcal{I}}}}}}\left({{\theta_{D}}}\right)$.
@@ IV-D Rank-one Constraint
The rank-one constraint is non-convex, which means the relevant toolbox can not directly solve the optimization problem. The above issues are solved by omitting the rank-one constraint and the rank of the obtained solution may be larger than one. Based on Refs. [18] and [24], the feasible suboptimal solution of the original problem is obtained by utilizing specific methods based on the SDR technology, and rank-one constraint is solved by Gaussian randomization, singular value decomposition, etc. It must be noted that these methods may result in a certain degree of distortion, which is tolerable in many cases. However, these methods cannot be utilized to solve the previously formulated problems since the recovered result may not satisfy ${\gamma_{D}}\leq{\gamma_{E}}$, which results in eavesdropping failure. In this work, an iterative method, SROCR, is utilized to solve this problem [42, 43]. Specifically, the rank-one constraint is relaxed by restricting the ratio of the maximum eigenvalue and trace of the matrix. In particular, for the scenario with ${\mathrm{rank}}\left({\mathbf{\tilde{R}}}\right)>1$, the eigenvalue decomposition of the matrix is firstly carried out and the eigenvalue matrix ${\mathbf{\tilde{\Sigma}}}={\mathrm{diag}}\left({{\lambda_{1}},...,{\lambda_{{N_{t}}}}}\right)$ is obtained. According to the relationship between the matrix eigenvalue and the matrix trace, the maximum eigenvalue of ${\mathbf{\tilde{R}}}$ must be less than the trace. The eigenvector of maximum eigenvalue of ${\mathbf{\tilde{R}}}$ is expressed as ${\mathbf{u}}$ and there must be ${{\mathbf{u}}^{H}}{\mathbf{\mathbf{\tilde{R}}}}{\mathbf{u}}<{\mathrm{tr}}\left({{\mathbf{\mathbf{\tilde{R}}}}}\right)$. By using the iterative method, the one constraint is transformed into
@@
Input: Initialize $P_{0}$, $\Delta\theta$, $\varphi$, $\gamma_{s}$, $\gamma_{D}$,$\gamma_{E}$, $\tau$, and $\vartheta$.
@@
The algorithm complexity analysis is as follows. $\mathcal{P}_{1.1}$ is composed of 1 linear matrix inequality (LMI) constraint of size $N_{t}$ and 2 LMI constraints of size 1, the computational complexity is $O\left({\sqrt{{N_{t}}+2}\left[{N_{t}^{2}\left({N_{t}^{3}+2}\right)+N_{t}^{4}\left({N_{t}^{2}+2}\right)+N_{t}^{6}}\right]}\right)$. $\mathcal{P}_{2.2}$ is composed by 1 LMI constraints of size $N_{t}$, ${2K{\varpi_{1}}+k{\varpi_{2}}+1}$ LMI constraints of size 1, the computational complexity is $O\left(\sqrt{{N_{t}}+\left({2K{\varpi_{1}}+K{\varpi_{2}}+1}\right)}\left[N_{t}^{2}\left({N_{t}^{3}+2K{\varpi_{1}}+K{\varpi_{2}}+1}\right)+\right.\right.$
$\left.\left.N_{t}^{4}\left({N_{t}^{2}+2K{\varpi_{1}}+K{\varpi_{2}}+1}\right)+N_{t}^{6}\right]\right)$, where ${\varpi_{1}}$ and ${\varpi_{2}}$ are cardinality of ${\Omega_{1}}$ and ${\Omega_{2}}$, respectively. $\mathcal{P}_{5.1}$ and $\mathcal{P}_{5.2}$ have 1 LMI constraints of size $N_{t}$, ${2K{\varpi_{1}}+k{\varpi_{2}}+2}$ LMI constraint of size 1, the computational complexity is $O\left(\sqrt{{N_{t}}+\left({2K{\varpi_{1}}+K{\varpi_{2}}+2}\right)}\left[N_{t}^{2}\left({N_{t}^{3}+2K{\varpi_{1}}+K{\varpi_{2}}+2}\right)+\right.\right.$
$\left.\left.N_{t}^{4}\left({N_{t}^{2}+2K{\varpi_{1}}+K{\varpi_{2}}+2}\right)+N_{t}^{6}\right]\right)$, and $\mathcal{P}_{\mathrm{5.{\mathrm{1b}}}}$ and $\mathcal{P}_{5.{\mathrm{2b}}}$ have 1 LMI constraints of size $N_{t}$, ${2K{\varpi_{1}}+k{\varpi_{2}}+3}$ LMI constraint of size 1, the computational complexity is $O\left(\sqrt{{N_{t}}+\left({2K{\varpi_{1}}+K{\varpi_{2}}+3}\right)}\left[N_{t}^{2}\left({N_{t}^{3}+2K{\varpi_{1}}+K{\varpi_{2}}+3}\right)+\right.\right.$
@@
$\left.\left.N_{t}^{4}\left({N_{t}^{2}+2K{\varpi_{1}}+K{\varpi_{2}}+3}\right)+N_{t}^{6}\right]\right)$. With the given convergence accuracy $\vartheta$, the computational complexity of algorithm 1 is expressed as $O\left({\sqrt{{N_{t}}+2K{\varpi_{1}}+K{\varpi_{2}}}N_{t}^{6}\log\left({\frac{1}{\vartheta}}\right)}\right)$ [19, 24], the computational complexity of algorithm 2 is $O\left({{I_{{\mathrm{iter}}}}}\left(\sqrt{{N_{t}}+2K{\varpi_{1}}+K{\varpi_{2}}}N_{t}^{6}\right)\log\left({\frac{1}{\vartheta}}\right)\right)$, while $I_{\mathrm{iter}}$ represents the number of iterations.
@@ V Numerical Results and Discussion
In this section, numerical results are presented to prove the convergence and effectiveness of our algorithm. The positive direction of the X axis is used as the reference, and the azimuth angles of $T_{k}$ $\left({K=2}\right)$ and $D$ relative to $E$ are expressed as ${\theta_{T_{k}}}$ and ${\theta_{D}}$, respectively. The direction of the radar detection range is $\left[{-{{90}^{\circ}},{{90}^{\circ}}}\right]$ , $\sigma^{2}$ is normalized to be unity [12], [38], and other parameters are listed in Table III. In all the figures, ‘PEC,’ ‘TSC,’ and ‘JPT’ correspond to the results for PE-centric, sensing-centric, and joint optimization of PE and TS, respectively.
TABLE III: Parameters Setting
@@
${d}_{ED}$ 1000 m $\theta_{T}$ $-25^{\circ}/25^{\circ}$
$\theta_{D}$ $60^{\circ}$ $\beta_{0}$ $-30$ dB
@@
$\sigma_{D}^{2}=\sigma_{E}^{2}$ $-80$ dBm $\varphi$ 0.05
@@
(b) CRB versus varying $P_{0}$.
(c) Eavesdropping rate versus varying $P_{0}$.
Figure 2: The results versus varying $P_{0}$ with $N_{t}=12$ and $\Delta\theta=5$.
@@
(b) Determinant of CRB versus varying $N_{t}$.
(c) Eavesdropping rate for varying $N_{t}$.
Figure 3: The results versus varying $N_{t}$ with $\rho=0.5$ and $\Delta\theta=5$.
@@
Fig. 2(a) plots the beampattern gains of the JPT scheme with varying power budgets. It can be observed that the gains of the mainlobe increases with increasing $P_{0}$, while the width of the mainlobes remains unchanged. Fig. 2(b) demonstrates the relationship between CRB and radar power budget for various schemes. It can be observed that the determinant of CRB decreases with the increase of $P_{0}$, which denotes that increasing the power budget can improve the accuracy of sensing. In the low-$P_{0}$ region, the TSC scheme’s sensing performance outperforms that of the JPT scheme, while in the high-$P_{0}$ region, the JPT scheme’s sensing performance outperforms that of the TSC scheme. This is because the beampattern gains in $D$’s direction is limited in the JPT scheme. With the increasing budget, more power is allocated to the mainlobe, enhancing the sensing performance. Fig. 2(c) plots the eavesdropping rate of various schemes versus the varying $P_{0}$. We can find that both the PEC and JPT ($\rho\neq 0$) schemes obtain the maximum eavesdropping rate while the JPT ($\rho=0$) scheme achieve slightly worse PE performance. The reason is $\rho=0$ in $\mathcal{P}_{5.{\mathrm{1b}}}$ denotes that the objection funciton is minimizing the CRB and only there is a minimum value for the gain in $D$’s direction given in (20d). Thus, a high power will result in a lower eavesdropping rate. The eavesdropping rate of the TSC scheme is worst and decreases with the increase of $P_{0}$ because (22e) and (22f) correspond to successful PE and the maximize interference at $D$, respectively. Obviously, larger $\alpha_{I}$, less eavesdropping rate due to $\alpha_{I}$ denotes the interference threshold at $D$.
@@
Fig. 3(a) shows the relationship between the beampattern gains of the JPT scheme and $N_{t}$. The results show that the mainlobe’s beampattern gains increases and its width decreases with increasing $N_{t}$. Fig. 3(b) demonstrates the sensing performance of all the schemes is improved with increasing $N_{t}$. Specifically, when the number of antennas is small, the TSC scheme’s sensing performance is better than that of the JPT scheme. As the number of antennas increases, the JPT scheme’s sensing performance increases faster than that of the TSC scheme. Moreover, the PEC scheme’s sensing performance tends to be constant, which outperforms that of the TSC scheme since the beampattern gains in $D$’s direction is optimized so that more power is allocated to the mainlobe in the PEC scheme. Fig. 3(c) plots the eavesdropping rate for all the schemes versus varying $N_{t}$. We can find that the PEC scheme achieves the maximum eavesdropping rate while the JPT scheme gradually reaches it as the number of antennas increases. The TSC scheme’s eavesdropping rate is the worst because only successful PE is guaranteed, and the minimum beampattern gains in the direction of $D$ is given. Moreover, the increasing of $N_{t}$ enhances the beam gain, so the interference in the $D$’s direction increases, and the eavesdropping rate decreases. An interesting result is that, for the TSC scheme, the eavesdropping rate with a higher power budget ($P_{0}=30$ dBm) underperforms that with a lower power budget ($P_{0}=27$ dBm). This is because there is a constraint on the minimum gain in $D$’s direction in (20d), and as the power budget increases, the gains (interference power) in $D$’s direction also increases.
@@
(b) Determinant of CRB with varying $\Delta\theta$.
(c) Eavesdropping rate for varying $\Delta\theta$.
@@
Fig. 4(a) plots the beampattern gains for the scenarios when target angles are perfectly known. It can be observed that the gain is maximum in the target direction and equal to the threshold in $D$’s direction. For the TSC scheme, the interference at $D$’s direction is slightly larger because the minimum gain in $D$’s direction is given in (20d). Fig. 4(b) demonstrates the CRB versus varying $\Delta\theta$ under different schemes. As the uncertain angle increases, the sensing performance of all the schemes deteriorates because the larger area needs to be covered with the given power. Moreover, the change of uncertain angle has a greater effect on the JPT’s sensing performance than that of the TSC scheme. The sensing performance of the JPT scheme is better than that of the TSC at scenarios with higher power, for the same reasons as Figs. 2(b) and 3(b). Fig. 4(c) shows the eavesdropping rate versus varying $\Delta\theta$ with different schemes. We can find that both the PEC and JPT schemes obtain the maximum eavesdropping rate, which is independent of the uncertain angle. For the TSC scheme, the eavesdropping rate increases with the increase of the uncertain angle because the range to be covered increases, and the interference in the $D$’s direction decreases. For a larger power ($P_{0}=30$ dBm), the interference in the $D$’s direction stays constant, thus the eavesdropping rate is independent of the uncertain angle.
@@
(b) Determinant of CRB with $P_{0}=26$ dBm.
(c) Determinant of CRB with $P_{0}=28$ dBm.
(d) Determinant of CRB with $P_{0}=30$ dBm.

### v3

@@ III-B Target Sensing Only
CRB is the lower bound of the unbiased estimation variance, and therefore minimizing CRB is a strategy for improving the accuracy of the estimation555 There are three different types of scalar CRB metrics: TCRB, MECRB, and DCRB [14, 41]. To comprehensively evaluate the joint estimation performance of multiple parameters, the DCRB should be prioritized. If a quick comparison of independent estimation accuracy for individual parameters is needed, the TCRB is the preferred choice. For ensuring controllable estimation error under worst-case conditions, the MECRB should be given priority. In this work, the DCRB is chosen as the performance metric of TS. The proposed schemes in this work also fit those scenarios wherein the trace and the maximum eigenvalue of the CRB are utilized as the matrices. . In this subsection, the transmitting beam is designed only to minimize the DCRB.
@@ III-B2 Imperfect Target Direction
In this subsection, a more practical scenario is considered wherein the estimated target angle $\theta_{T_{k}}$} is imperfect and there is an uncertain angle $\Delta\theta$. Specifically, the angle of the $k$-th target varies in ${\Omega_{1}}=\left[{{\theta_{T_{k}}}-\Delta\theta,{\theta_{T_{k}}}+\Delta\theta}\right]$. Then, the following optimization problem is formulated
@@ IV Jointly Optimization of Proactive Eavesdropping and TS
where $\alpha_{\Phi}$ denotes the target threshold for DCRB.
@@ IV-A Proactive Eavesdropping-centric Design
where (20e) is the TS constraint. $\mathcal{P}_{3}$ is a convex problem after dropping constraint (20c), which can be solved by CVX toolbox.
@@ IV-B Target Sensing-centric Design
where $\alpha_{I}$ denotes the interference threshold at $D$. It must be noted that (21f) and (22f) denote the requirement for the ER. Obviously, $\alpha_{I}$ has an essential effect on the optimized results. $\mathcal{P}_{4.1}$ and $\mathcal{P}_{4.2}$ can be transformed into a convex problem by means of SDR and are solved directly using the CVX toolbox when (21c) and (22c) are omitted.
@@
In this subsection, a normalized weighted optimization problem that simultaneously considers both PE and TS performance is proposed. Design of the weighted optimization between the DCRB and ER presents the challenge that both performance metrics have different units and potentially different magnitudes. Inspired by [26], both $\mathcal{I}\left({{\theta_{D}}}\right)$ and $\left|{{\bm{\Phi}}}\left({{\bm{\zeta}}}\right)\right|$} are normalized with their respective bound obtained in $\mathcal{P}_{1.1}$ and $\mathcal{P}_{2.1}$, respectively.
@@ IV-C1 Joint Optimization Proactive Eavesdropping and Target Sensing
To simultaneously consider PE and TS, while taking the uncertain angle of $T_{k}$ into account, the weighted optimization problem is formulated as666 It should be noted that (23h) and (23i) correspond to $\rho=1$ and $\rho=0$, respectively. In particular, $\mathcal{P}_{5.1}$ degenerates into $\mathcal{P}_{3}$ and $\mathcal{P}_{4.2}$ when $\rho=1$ and $\rho=0$, respectively.
@@
${{\left|{{\bm{\Phi}}}\left({{\bm{\zeta}}}\right)\right|}\leq\alpha_{\Phi}},$ (23h)
${{\mathcal{I}\left({{\theta_{D}}}\right)\leq\alpha_{I}}},$ (23i)
where $\mathcal$$\tilde${I}} and $\tilde{\bm{\Phi}}$}}} are the results of $\mathcal{P}_{1.1}$ and $\mathcal{P}_{2.1}$ $\left(\mathcal{P}_{2.2}\right)$, respectively, and $0\leq\rho\leq 1$ signifies the weights factor between PE and TS. $\mathcal{P}_{5.1}$ is an SDR problem by neglecting (23e) and can be solved with the CVX toolbox.
@@ IV-C2 The interference power is zero
It is unnecessary for $E$ to send AN to $D$ when the channel quality of SES-E is better than that of SDS-D, i.e. $\frac{{h_{SD}^{2}}}{{\sigma_{D}^{2}}}\leq\frac{{h_{SE}^{2}}}{{\sigma_{E}^{2}}}$ in (23d) [44]. In these scenarios, the ER is equal to the receiving rate of $D$ without interference and $\mathcal{P}_{3}$ is transformed into a TS sub-problem, which is expressed as
@@
   Solve $\mathcal{P}_{1.1}$ for given parameters and obtain the solution ${{{\mathrm{\tilde{\mathcal{I}}}}}}\left({{\theta_{D}}}\right)$.
@@ IV-D Rank-one Constraint
[…]
@@
${{\left|{{\bm{\Phi}}}\left({{\bm{\zeta}}}\right)\right|}\leq\alpha_{\Phi}},$ (26h)
${{\mathcal{I}\left({{\theta_{D}}}\right)\leq\alpha_{I}}},$ (26i)
@@
Input: Initialize $P_{0}$, $\Delta\theta$, $\varphi$, $\gamma_{s}$, $\gamma_{D}$$\gamma_{E}$, $\tau$, and $\vartheta$.
@@
IV-E Comparison of formulated problems
In this work, there are eight (two kinds of) formulated problems, which are introduced as follows.
  • The first kind of problem is considering PE or TS, respectively.
    1. 1.
      In $\mathcal{P}_{1}$ ($\mathcal{P}_{1.1}$), the transmitting beam is designed only to maximize the ER, considering successful PE, and the TS is not considered.
    2. 2.
      In $\mathcal{P}_{2.1}$ and $\mathcal{P}_{2.2}$, the transmitting beam is designed only to minimize the DCRB and the PE is not considered. $\mathcal{P}_{2.1}$ assumes that the azimuth angles of targets ($\theta_{T_{k}}$}) are perfect, while $\mathcal{P}_{2.2}$ examines the imperfect $\theta_{T_{k}}$}.
    3. 3.
      The solutions of $\mathcal{P}_{1}$ and $\mathcal{P}_{2.2}$ are utilized in $\mathcal{P}_{5.1}$ and $\mathcal{P}_{5.2}$.
  • The second kind of problem is considering PE and TS simultaneously.
    1. 1.
      In $\mathcal{P}_{3}$, the transmit covariance matrix is optimized to maximize the PE performance (minimize the beampattern gains at $D$) while taking the TS constraint into account, which is called PE-centric (PEC) design.
    2. 2.
      In $\mathcal{P}_{4.1}$ and $\mathcal{P}_{4.2}$, the transmit covariance matrix is optimized to maximize the TS performance (minimize the estimation DCRB) while ensuring successful PE and ER constraints, which is called the TS-centric (TSC) design. $\mathcal{P}_{4.1}$ assumes that the azimuth angles of targets ($\theta_{T_{k}}$}) are perfect, while $\mathcal{P}_{4.2}$ examines the imperfect $\theta_{T_{k}}$}.
    3. 3.
      In $\mathcal{P}_{5.1}$ and $\mathcal{P}_{5.2}$, the normalized weighted summation between $\left|{{\bm{\Phi}}}\left({{\bm{\zeta}}}\right)\right|$} and $\mathcal{I}\left({{\theta_{D}}}\right)$ is minimized considering successfully PE, DCRB requirements, and interference requirement at $D$, which is called as joint optimization of PE and TS (JPT) design. Different from $\mathcal{P}_{5.1}$, $\mathcal{P}_{5.2}$ considers a special scenario wherein the channel quality of SES-E is better than that of SDS-D, i.e. $\frac{{h_{SD}^{2}}}{{\sigma_{D}^{2}}}\leq\frac{{h_{SE}^{2}}}{{\sigma_{E}^{2}}}$. In this case, $E$ does not need to send AN to $D$, and all the power can be utilized to TS.
Table III summarizes the difference for all the formulated problems.
TABLE III: Comparison of formulated problems.
Problem
PE
TS
Perfect/imperfect
Target Direction
$\frac{{h_{SD}^{2}}}{{\sigma_{D}^{2}}}\leq\frac{{h_{SE}^{2}}}{{\sigma_{E}^{2}}}$
$\mathcal{P}_{1}$ ($\mathcal{P}_{1.1}$)
$\mathcal{P}_{2.1}$ Perfect
$\mathcal{P}_{2.2}$ Imperfect
$\mathcal{P}_{3}$ Constraint
$\mathcal{P}_{4.1}$ Constraint Perfect
$\mathcal{P}_{4.2}$ Constraint Imperfect
$\mathcal{P}_{5.1}$ ($\mathcal{P}_{5.{\mathrm{1b}}}$) Imperfect
$\mathcal{P}_{5.2}$ ($\mathcal{P}_{5.{\mathrm{2b}}}$) Imperfect
IV-F Convergence Discussion and Complexity Analysis
While optimization problems involving multi-variable coupling often necessitate iterative approaches such as BCD or SCA. The convergence analysis serves as key theoretical validation. This work formulates problems containing only a single variable, $\mathbf{R}$. By applying SDR to relax the rank-one constraint, the problem is transformed into a convex optimization task that can be solved directly. Although this relaxation reduces computational complexity, it may introduce a performance gap between the relaxed solution and the true optimum of the original non-convex problem. To mitigate this gap and enhance solution tightness, we utilize an iterative optimization method based on rank-one relaxation. Specifically, we define the ratio of the maximum eigenvalue to the matrix trace as a relaxation performance metric and adopt a binary search strategy to refine the solution while ensuring constraint feasibility iteratively. Crucially, this approach guarantees strict improvement in the relaxation metric with each iteration. By setting a convergence threshold (e.g., 0.999), the final solution’s performance index is driven arbitrarily close to 1, achieving a high-precision approximation of the original problem’s optimum. This method not only circumvents the computational challenges of directly solving non-convex problems but also significantly improves accuracy through iterative refinement, offering a practical and theoretically grounded solution pathway for rank-one-constrained non-convex optimization.
[…]
@@ IV-F Convergence Discussion and Complexity Analysis
$\left.\left.N_{t}^{4}\left({N_{t}^{2}+2K{\varpi_{1}}+K{\varpi_{2}}+1}\right)+N_{t}^{6}\right]\right)$, where $\varpi_{1}$} and $\varpi_{2}$} are cardinality of $\Omega_{1}$} and $\Omega_{2}$}, respectively. $\mathcal{P}_{5.1}$ have 1 LMI constraints of size $N_{t}$, $2K{\varpi_{1}}+K{\varpi_{2}}+4$} LMI constraint of size 1, $\mathcal{P}_{5.2}$ have 1 LMI constraints of size $N_{t}$, $2K{\varpi_{1}}+K{\varpi_{2}}+2$} LMI constraint of size 1, $\mathcal{P}_{\mathrm{5.{\mathrm{1b}}}}$ have 1 LMI constraints of size $N_{t}$, $2K{\varpi_{1}}+K{\varpi_{2}}+5$} LMI constraint of size 1, $\mathcal{P}_{5.{\mathrm{2b}}}$ have 1 LMI constraints of size $N_{t}$, $2K{\varpi_{1}}+K{\varpi_{2}}+3$} LMI constraint of size 1. With the given convergence accuracy $\vartheta$, the computational complexity of algorithm 1 is expressed as $O\left({\sqrt{{N_{t}}+2K{\varpi_{1}}+K{\varpi_{2}}}N_{t}^{6}\log\left({\frac{1}{\vartheta}}\right)}\right)$ [21, 26], the computational complexity of algorithm 2 is $O\left({{I_{{\mathrm{iter}}}}}\left(\sqrt{{N_{t}}+2K{\varpi_{1}}+K{\varpi_{2}}}N_{t}^{6}\right)\log\left({\frac{1}{\vartheta}}\right)\right)$, while $I_{\mathrm{iter}}$ represents the number of iterations.
@@ V Numerical Results and Discussion
In this section, numerical results are presented to prove the convergence and effectiveness of our algorithm. The positive direction of the X axis is used as the reference, and the azimuth angles of $T_{k}$ $\left({K=2}\right)$ and $D$ relative to $E$ are expressed as $\theta_{T_{k}}$} and $\theta_{D}$}, respectively. The direction of the radar detection range is $\left[{-{{90}^{\circ}},{{90}^{\circ}}}\right]$ , $\sigma^{2}$ is normalized to be unity [14], [41], and other parameters are listed in Table IV.
TABLE IV: Parameters Setting
@@
${d}_{ED}$ 1000 m $\theta_{T}$ $25^{\circ}/25^{\circ}$
$\theta_{D}$ $60^{\circ}$ $\beta_{0}$ 30-30 dB
@@
$\sigma_{D}^{2}=\sigma_{E}^{2}$ 80-80 dBm $\varphi$ 0.05
@@
(b) DCRB versus varying $P_{0}$.
(c) ER versus varying $P_{0}$.
Figure 2: The results versus varying $P_{0}$ with $N_{t}=12$ and $\Delta\theta=3$.
@@
(b) DCRB versus varying $N_{t}$.
(c) ER for varying $N_{t}$.
Figure 3: The results versus varying $N_{t}$ with $\rho=0.5$ and $\Delta\theta=3$.
@@
Fig. 2(a) plots the beampattern gains of the JPT scheme with varying power budgets. It can be observed that the gains of the mainlobe increases with increasing $P_{0}$, while the width of the mainlobes remains unchanged. Fig. 2(b) demonstrates the relationship between DCRB and radar power budget for various schemes. It can be observed that the DCRB decreases with the increase of $P_{0}$, which denotes that increasing the power budget can improve the accuracy of sensing. When $\rho=0$, JPT is equivalent to the TSC scheme; when $\rho=1$, JPT is equivalent to the PEC scheme. The JPT scheme’s sensing performance outperforms that of the TSC scheme. This is because the beampattern gains in $D$’s direction is limited in the JPT scheme. With the increasing budget, more power is allocated to the mainlobe, enhancing the sensing performance. Fig. 2(c) plots the ER of various schemes versus the varying $P_{0}$. We can find that both the PEC and JPT ($\rho\neq 0$) schemes obtain the maximum ER. The ER of the TSC and JPT with $\rho=0$ are the worst exhibits no significant regular variation with the increase of $P_{0}$. That is because (22e) and (22f) correspond to successful PE and the maximum interference at $D$, respectively and no maximizing ER in TSC and JPT with $\rho=0$. Obviously, larger $\alpha_{I}$, less ER due to $\alpha_{I}$ denotes the interference threshold at $D$.
@@
Fig. 3(a) shows the relationship between the beampattern gains of the JPT scheme and $N_{t}$. The results show that the mainlobe’s beampattern gains increases and its width decreases with increasing $N_{t}$. Fig. 3(b) demonstrates the sensing performance of all the schemes is improved with increasing $N_{t}$. Specifically, when the number of antennas is small, the TSC scheme’s sensing performance is better than that of the JPT scheme. As the number of antennas increases, the JPT scheme’s sensing performance increases faster than that of the TSC scheme. Moreover, the PEC scheme’s sensing performance tends to be constant, which outperforms that of the TSC scheme since the beampattern gains in $D$’s direction is optimized so that more power is allocated to the mainlobe in the PEC scheme. Fig. 3(c) plots the ER for all the schemes versus varying $N_{t}$. We can find that both the PEC and JPT schemes reach the maximum ER at the beginning, and this rate remains constant regardless of the number of antennas. The TSC scheme’s ER is the worst because only successful PE is guaranteed, and the minimum beampattern gains in the direction of $D$ is given. Moreover, the increase of $N_{t}$ enhances the beam gain, so the interference in the $D$’s direction increases, and the ER decreases. An interesting result is that, for the TSC scheme, the ER with a higher power budget ($P_{0}=30$ dBm) underperforms that with a lower power budget ($P_{0}=28$ dBm). This is because there is a constraint on the minimum gain in $D$’s direction in (20d), and as the power budget increases, the gains (interference power) in $D$’s direction also increases.
@@
(b) DCRB with varying $\Delta\theta$.
(c) ER for varying $\Delta\theta$.
@@
Fig. 4(a) plots the beampattern gains for the scenarios when target angles are perfectly known. It can be observed that the gain is maximum in the target direction and equal to the threshold in $D$’s direction. For the TSC scheme, the interference at $D$’s direction is slightly larger because the minimum gain in $D$’s direction is given in (20d). The essence of JPT design lies in the directional redistribution of beam energy by optimizing the gain of the target-direction beam while constraining energy allocation in specific directions (e.g., $D$-direction). This mechanism directly leads to its differentiated performance in non-target directions compared to TSC and PEC schemes, further validating the effectiveness of resource-coupled optimization in ISAC systems. Fig. 4(b) demonstrates the DCRB versus varying $\Delta\theta$ under different schemes. As the uncertain angle increases, the sensing performance of all the schemes deteriorates because the larger area needs to be covered with the given power. Moreover, the change of uncertain angle has a greater effect on the JPT’s sensing performance than that of the TSC scheme. The sensing performance of the JPT scheme is better than that of the TSC at scenarios with higher power, for the same reasons as Figs. 2(b) and 3(b). Fig. 4(c) shows the ER versus varying $\Delta\theta$ with different schemes. We can find that both the PEC and JPT schemes obtain the maximum ER, which is independent of the uncertain angle. For the TSC scheme, the ER increases with the increase of the uncertain angle because the range to be covered increases, and the interference in the $D$’s direction decreases. For a larger power ($P_{0}=32$ dBm), the interference in the $D$’s direction stays constant, thus the ER is independent of the uncertain angle.
@@
(b) DCRB with $P_{0}=28$ dBm.
(c) DCRB with $P_{0}=30$ dBm.
(d) DCRB with $P_{0}=32$ dBm.